OpenStudy (anonymous):

What is the solution to the rational equation (4x/3)+(1/3) = 5 ?

OpenStudy (raden):

times 3 from both sides, what u get ...

OpenStudy (anonymous):

i don't get it :(

OpenStudy (anonymous):

4x/3 + 1/3 =5 4x+1=15 4x=14 x=14/4=3.5

OpenStudy (anonymous):

thank you so much i get now